# Dispersion on Time-Varying Graphs

### v1

@@ 3.1 Algorithm for general graph
        Node $\bm{u}$ has no agent with $\bm{setlled=1}$: If $a_{i}$ is the minimum ID at node $u$, it settles at node $u$ and updates $a_{i}.settled=1$, $a_{i}.grp\_label=\bot$ and $a_{i}^{u}(G_{2})$ to $(a_{i}.prt\_in,a_{i}.grp\_label,a_{i}.dfs\_label)$. If there is any unsettled agent $a_{j}$ from $G_{1}$ at node $u$, then $a_{i}$ updates $a_{i}^{u}(G_{1})$ to $(a_{j}.prt\_in,\\ a_{j}.grp\_label,a_{j}.dfs\_label)$. If $a_{i}$ is not the minimum ID, then it updates $a_{i}.prt\_out=(a_{i}.prt\_in+1)$ mod $\delta_{u}$. If $a_{i}.prt\_out\neq a_{i}.prt\_in$, then it moves through $a_{i}.prt\_out$. Else, if $a_{i}.prt\_out=a_{i}.prt\_in$, then it sets $a_{i}.state=backtrack$ and move through $a_{i}.prt\_out$.
@@
The pseudocode of our algorithm, Dispersion in 1-TVG with $n+1$ agents, is provided in Algorithm 1.
@@ 56                               It does not do anything
3.2 Correctness and analysis of algorithm
Proof
Proof
Based on Claim 1 &\&& 2, we have the following lemma.
Lemma 1
Our algorithm solves Dispersion on TVG for $k=n+1$ agents correctly.
Proof
Due to Claim 1 &\&& 2, we can say either $G_{1}$ or $G_{2}$ is dispersed and unsettled agents in $G_{1}\,(G_{2})$ understand whether agents in $G_{2}\,(G_{1})$ are dispersed. And, agents in $G_{1}\,($or $G_{2})$ divide into two new groups. In this way, in the end, group $G_{1}$ (and $G_{2}$) contains 1 agent. Both groups start running our algorithm, due to Claim 1 &\&& 2, we can say that either agent in $G_{1}$ or $G_{2}$ is settled at some node. Therefore, our algorithm solves Dispersion on TVG for $k=n+1$ correctly. ∎
Lemma 2
The time complexity of our algorithm is $O(n^{4}\cdot\log n)$.
Proof
Using Lemma 1, if $G_{1}$ (or $G_{2}$) understands that $G_{2}$ (or $G_{1}$) is dispersed, then $G_{1}$ (or $G_{2}$) divides into two new groups, and both groups contain half agents of $G_{1}$ (or $G_{2}$). These two groups repeat the same procedure. The number of such group divisions is $O(\log n)$, and one of two groups is dispersed in $64n^{4}+8m$ rounds. Therefore, our algorithm takes $O(n^{4}\cdot\log n)$ rounds to solve Dispersion on TVG for $k=n+1$. ∎
The following lemma is w.r.t. memory lower bound. The proof of this lemma is motivated from [2]. In [2], they have provided the memory lower bound to solve dispersion on a static graph when $k=n$. We extend their idea for the memory lower bound when $k>n$.
Lemma 3
Assuming all agents are given the same amount of memory, agents require $\Omega(\log n)$ bits of memory each for any deterministic algorithm to solve Dispersion on TVG when $k\geq n+1$.
Proof
Suppose all agents have $o(\log n)$ bits of memory. Each agent’s state space is $2^{o(\log n)}=n^{o(1)}$. Since there are at least $n+1$ agents, by pigeonhole principle, there exist at least $k-n+2$ agents with the same state space. Let agents $a_{1}$, $a_{2}$, …, $a_{k-n+2}$ have the same state space. Let us suppose that all agents are initially co-located at some node. Since all agents run the same deterministic algorithm, and agents $a_{1}$, $a_{2}$, …, $a_{k-n+2}$ are co-located initially, they will run the same moves. It means that agents $a_{1}$, $a_{2}$, …, $a_{k-n+2}$ remain stay together. Therefore, by the end of the algorithm, agents $a_{1}$, $a_{2}$, …, $a_{k-n+2}$ are at the same node. Solving Dispersion on TVG requires a configuration where there is at least one agent at every node. We will never solve Dispersion on TVG because there is no way for agents $a_{1}$, $a_{2}$, …, $a_{k-n+2}$ to settle down on at least two different nodes. Therefore, all $k$ agents are at most $n-1$ nodes by the end of the algorithm. Thus, to solve Dispersion on TVG, agents need $\Omega(\log n)$ memory each. ∎
Lemma 4
In our algorithm, agents require $\Theta(\log n)$ memory.
Proof
Now we are ready to provide our main theorem  [^1] Our algorithm can be extended to solve Dispersion on TVG when $k>n$. Since agents are co-located they can consider $k$ as an upper bound of $n$. In our algorithm in place of $n$, if we use $k$, our algorithm solves Dispersion on TVG with $k>n+1$ co-located agents and the time complexity is $O(k^{4}\cdot\log n)$..
Theorem 3.1
Our algorithm solves Dispersion on TVG for $k=n+1$ in $\mathcal{G}$ in $O(n^{4}\cdot\log n)$ rounds using $\Theta(\log n)$ memory per agent.
Proof
The proof follows from Lemma 1, 2, and 4. ∎
Observation 1
All agents know the value of $n$, and $m\leq n^{2}$. If all agents want to achieve the termination, can calculate the value of $72n^{4}\cdot\lceil\log n\rceil$ and terminate after $72n^{4}\cdot\lceil\log n\rceil$ rounds.
4 Conclusion

### v2

@@ 3.1 Algorithm for general graph
        Node $\bm{u}$ has no agent with $\bm{setlled=1}$: If $a_{i}$ is the minimum ID at node $u$, it settles at node $u$ and updates $a_{i}.settled=1$, $a_{i}.grp\_label=\bot$ and $a_{i}^{u}(G_{2})$ to $(a_{i}.prt\_in,a_{i}.grp\_label,a_{i}.dfs\_label)$. If there is any unsettled agent $a_{j}$ from $G_{1}$ at node $u$, then $a_{i}$ updates $a_{i}^{u}(G_{1})$ to $(a_{j}.prt\_in,a_{j}.grp\_\\ label,a_{j}.dfs\_label)$. If $a_{i}$ is not the minimum ID, then it updates $a_{i}.prt\_out=(a_{i}.prt\_in+1)$ mod $\delta_{u}$. If $a_{i}.prt\_out\neq a_{i}.prt\_in$, then it moves through $a_{i}.prt\_out$. Else, if $a_{i}.prt\_out=a_{i}.prt\_in$, then it sets $a_{i}.state=backtrack$ and move through $a_{i}.prt\_out$.
@@
The pseudocode of our algorithm is provided in Section 6.
3.2 Correctness and analysis of algorithm
Proof
Proof
Based on Claim 1 &\&& 2, we have the following lemma.
Lemma 2
Our algorithm correctly solves $(n+1)$-Dispersion on TVG for co-located $n+1$ agents.
Proof
Due to Claim 1 &\&& 2, we can say either $G_{1}$ or $G_{2}$ is dispersed and unsettled agents in $G_{1}\,(G_{2})$ understand whether agents in $G_{2}\,(G_{1})$ are dispersed. And, agents in $G_{1}\,($or $G_{2})$ divide into two new groups. In this way, in the end, group $G_{1}$ (and $G_{2}$) contains 1 agent. Both groups start running our algorithm, and due to Claim 1 &\&& 2, we can say that either agent in $G_{1}$ or $G_{2}$ is settled at some node. Therefore, our algorithm correctly solves $(n+1)$-Dispersion on TVG for co-located $n+1$ agents. ∎
Lemma 3
The time complexity of our algorithm is $O(n^{4}\cdot\log n)$.
Proof
Using Lemma 2, if $G_{1}$ (or $G_{2}$) understands that $G_{2}$ (or $G_{1}$) is dispersed, then $G_{1}$ (or $G_{2}$) divides into two new groups, and both groups contain half agents of $G_{1}$ (or $G_{2}$). These two groups repeat the same procedure. The number of such group divisions is $O(\log n)$, and one of two groups is dispersed in $64n^{4}+8m$ rounds. Therefore, our algorithm takes $O(n^{4}\cdot\log n)$ rounds to solve $(n+1)$-Dispersion on TVG for co-located $n+1$ agents. ∎
Lemma 4
In our algorithm, agents require $\Theta(\log n)$ memory.
Proof
Now, we are ready to provide our main theorem.
Theorem 3.1
Our algorithm solves $k$-Dispersion on TVG for $k=n+1$ co-located agents in $\mathcal{G}$ in $O(n^{4}\cdot\log n)$ rounds using $\Theta(\log n)$ memory per agent.
Proof
The proof follows from Lemma 2, 3, and 4. ∎
Observation 1
All agents know the value of $n$ and $m\leq n^{2}$. If all agents want to achieve the termination, can calculate the value of $72n^{4}\cdot\lceil\log n\rceil$ and terminate after $72n^{4}\cdot\lceil\log n\rceil$ rounds.
4 Discussion
In this section, we give a discussion of how to solve $(pn+1)$-Dispersion on TVG without global communication using 0-hop visibility. We also provide an algorithm to solve $(pn)$-Dispersion on TVG with global communication and 1-hop visibility using the algorithm of [16].
4.1 $(pn+1)$-Dispersion on TVG
In this section, we demonstrate how to extend the algorithm from Section 3 to address the $(pn+1)$-Dispersion on TVGs, where $p>1$. In the dispersed configuration, each node contains $p$ agents, except one node contains $p+1$ agents. We assume that agents are aware of the parameter $n$, and $pn+1$ agents are co-located at some node $v$. In the first round, agents distribute into $n$ groups, say $g_{1}$, $g_{2}$, …, $g_{n}$. Since agents are co-located, they can compute the value of $p$ using $n$. The number of agents in each $g_{i}$, $11\leq i\leq n-$ 1, is $p$, and $g_{n}$ is $p+1$. Each agent $a_{i}$ in $g_{j}$ remembers $j$ in $a_{i}.ID_{1}$ parameter, where $1\leq j\leq n$. In the same manner, agents start executing the algorithm mentioned in Section 3. In this algorithm, whenever agents find a node without any agents, then $g_{i}$ gets settled in place of an agent. Here, the remaining unsettled $g_{i}$s divide into two groups, say $G_{1}$ and $G_{2}$, in place of dividing groups of agents into 2 groups like Section 3. In Lemma 2, we have shown one of $G_{1}$ or $G_{2}$ gets dispersed based on their parameters. When $G_{1}$ ($G_{2}$) is left with one $g_{j}$ (unsettled group), and $G_{1}$ ($G_{2}$) is sure that $G_{2}$ ($G_{1}$) has been dispersed, then they follow the following steps. W.l.o.g., let $G_{1}$ be with one $g_{i}$ at node $v$, and $G_{2}$ has been dispersed. In this case, it is important to note that 1 node is left, which has no agents. If $v$ is such a node, then dispersion has been achieved. If $v$ is not such a node, then there is some node $u$ which has no agent. In our algorithm mentioned in Section 3, there were two agents left at the end who were searching for the single empty node via different paths, one gets to that node as the adversary can not stop both at a round, and the other settled when it is sure that it the only one that is left to settle. Here, we can not use our algorithm idea because the adversary can always stop one group, i.e., the last remaining group, to find the single empty node in the graph. Below we provide the idea with correctness how the group $g_{i}$ gets settled at node $u$. There are two cases, and we discuss each case as follows.
  1. 1.
    Case $\bm{|g_{i}|=p+1}$ : The agents start running the algorithm mentioned in Section 3 in the following manner. Now $g_{i}$ executes DFS for the single node, say $u$, that has less than $p$ agents, actually no agents at this point of time, on it. Whenever the missing edge appears, $g_{i}$ divides into two groups, say $G_{1}$ and $G_{2}$, and starts executing DFS. As per Lemma 2, one of the groups $G_{1}$ or $G_{2}$ finds the node $u$ that has $<p$ agents and settles there. This continues until $u$ has at least $p$ settled agents. In the end, like our algorithm, when $G_{1}$ ($G_{2}$) is left with one agent, and it is sure that $G_{2}$ ($G_{1}$) has been dispersed, it settles at the current node, say $v$, as by this time the remaining $p$ agents of $g_{i}$ must have settled at $u$. Note that $v$ can be the node $u$ as well, but irrespective of what $v$ is, the dispersed configuration is achieved as exactly one node ($v$) contains $p+1$ agents.
  2. 2.
    Case $\bm{|g_{i}|=p}$ : In this case, $g_{j}$ with $p+1$ has been settled at some node $u_{1}$ in the earlier round. In observation 1, we have given the number of rounds of termination. In this case, agents in $g_{i}$ wait till that round. Agents in $g_{i}$ become group $G_{1}$, and the largest ID at $g_{j}$ becomes $G_{2}$. They start executing DFS as per our algorithm. As per Claim 1, if $G_{1}$ waits for consecutive $16n^{2}$ odd rounds, then $G_{2}$ visits each node at least once. In this case, $G_{2}$ and $G_{1}$ meet, and they become a group of $p+1$ agents, which is an earlier case. If $G_{1}$ does not wait for consecutive $16n^{2}$ odd rounds, then it is moving on the current DFS. And in at most $16n^{2}\times 4n^{2}$ odd rounds, it reaches node $u$ and settles. In at most $64n^{2}$ odd rounds, the agent in $G_{2}$ becomes settled at its current node. This ensures that the agents achieve a dispersed configuration because, after $64n^{2}$ odd rounds, if $G_{2}$ has not met $G_{1}$, it indicates that $G_{1}$ has completed its depth-first search (DFS).
Time Complexity of solving $(pn+1)$-Dispersion on TVG:
Now we are ready to state the final theorem.
Theorem 4.1
$(pn+1)$-Dispersion on TVG can be solved in $O(n^{4}\cdot max(\log n,\,\log p))$ rounds using $\Theta(\log n)$ memory per agent.
Proof
The correctness is justified in the above discussion. Since agents are using the algorithm mentioned in Section 3. Therefore, $\Theta(\log n)$ memory per agent is required. Since $n$ many $g_{i}$s are executing the algorithm, therefore, in $O(n^{4}\cdot\log n)$ rounds, one of the aforementioned cases is possible. In the first case, agents take $O(n^{4}\cdot\log p)$ extra rounds to achieve the dispersed configuration, and in the second case if both groups meet with each other within $64n^{2}$ odd rounds, then they convert into the first case, and take $O(n^{4}\cdot\log p)$ extra rounds to achieve the dispersed configuration. In the second case, if they do not meet with each other, then $g_{i}$ has reached node $u$ within $64n^{2}$, and the dispersion has been achieved. Therefore, the total time is $O(n^{4}\cdot\log n)+O(n^{4}\cdot\log p)=O(n^{4}\cdot max(\log n,\,\log p))$. This completes the proof. ∎
4.2 $k$-Dispersion on TVG
In Section 2, we have shown that to solve $(pn)$-Dispersion on TVG, agents need 1-hop visibility and global communication. In this section, we will extend the algorithm of [16] to solve $k$-Dispersion on TVG when agents are equipped with 1-hop visibility and global communication. Recall that their algorithm works in the weak dynamic graph model in comparison to TVG, i.e., the adversary can delete or add edges arbitrarily, maintaining the graph in each round connected. Let’s refer to the algorithm as $Algo\_weak\_Disp(k)$. First, we recall a few results from their work.
Theorem 4.2
[16] Given $k\leq n$ agents placed arbitrarily on the nodes of any $n$-node graph $\mathcal{G}_{r}$, where $\mathcal{G}_{r}$ is the graph at round $r\geq 0$. $Algo\_weak\_Disp(k)$ solves the dispersion in $\Theta(k)$ rounds with $\Theta(\log k)$ bits at each agent in the synchronous setting with global communication and 1-hop visibility. Also, all agents understand that dispersion has been achieved and terminated.
Based on $Algo\_weak\_Disp(k)$, the authors have the following lemma.
Lemma 5
[16] Consider any $n$-node dynamic graph $\mathcal{G}_{r}$ at round $r\geq 0$. If $k\leq n$ agents are positioned on $l$ nodes of $\mathcal{G}_{r}$ at the beginning of round $r$, then at the beginning of round $r+1$, the agents are positioned on at least $l+1$ nodes of $\mathcal{G}_{r}$.
Now, we are ready to provide how we can extend $Algo\_weak\_Disp(k)$ to solve $k$-Dispersion on TVG when $k$ agents in $G$ need 1-hop visibility and global communication. Note that this idea works even if agents are scattered, agents do not know $n$, and the adversary deletes more than 1 edge while maintaining $\mathcal{G}$ connected.
The agents run $Algo\_weak\_Disp(k)$. Using Lemma 5, we can observe that the number of nodes without any agents is reduced by at least 1 in each round. Therefore, after $n$ rounds, all nodes have at least 1 agent. At each round $r$, agents check whether their neighbour is occupied by some agent or not and share this information using global communication with all agents. At some round $r$, when agents do not get such information, agents can understand that each node has at least 1 agent. At every round $r^{\prime}>r$, at each node $v$, the largest ID agent becomes a leader, and it shares the information of the number of agents at node $v$ along with its ID. In this way, at every round $r^{\prime}$ with the help of global communication, agents can learn the value of $k$ using the number of agents at other nodes, and they can also learn the value of $n$ based on the number of distinct IDs they receive. Therefore, they do not have to remember $k$ and $n$, and in each round $r^{\prime}$, they can learn $k$ and $n$. This information allows agents to compute $\lceil k/n\rceil$. From round $r+2$, they start executing $Algo\_weak\_Disp(k)$. In each round, agents can learn $\lceil k/n\rceil$. This is important otherwise the memory requirement per agent will be $O(max(\log p,\log n))$. When they find at least one node has less than $\lceil k/n\rceil$ agents. They can follow $Algo\_weak\_Disp(k)$. Using Lemma 5, we can say that if there is a node $u$ with more than $\lceil k/n\rceil$ agents, then the extra agent from node $u$ moves to the other node $v$ which has less than $\lceil k/n\rceil$ agents. In this way, they can do this in each round. When agents find no node with more than $\lceil k/n\rceil$ agents, they can understand $k$-Dispersion on TVG has been achieved. Since $k$ agents are present, agents can achieve $k$-Dispersion on TVG in $O(k)$ rounds. In each round, agents need to remember their ID, and it does not have to remember $k$ or $n$. Therefore, agents need $O(\log n)$ memory as ID range is $[1,\,n^{c}]$, where $c$ is some constant. Using Lemma 1, we can say $\Theta(\log n)$ memory per agent is required. Based on this, we have the final theorem.
Theorem 4.3
Let $k$ agents be placed arbitrarily on the nodes of $G$, and the adversary can delete edges arbitrarily from footprint $G$, maintaining $\mathcal{G}$ connected. $Algo\_weak\_Disp(k)$ solves $k$-Dispersion on TVG in $O(k)$ rounds with $\Theta(\log n)$ bits of memory at each agent with global communication and 1-hop visibility. Also, all agents understand that dispersion has been achieved and terminated.
5 Conclusion
6 Pseudocodes